OpenStudy (anonymous):

factor using zero principle (x+6)(x-2)=9

OpenStudy (anonymous):

x^2 + 4x -12 = 9 x^2 + 4x -21 = 0 (x - 3)(x + 7) = 0 x = 3 or x = -7